OpenStudy (anonymous):

Solve by substitute {2x-3y=8 {x+y=11 Please help

OpenStudy (anonymous):

so x= 11-y now plug that into the first equation: 2(11-y) -3y = 8 now solve that

OpenStudy (anonymous):

once you get the answer for that it will be y = ... and then you plug that answer back into the second equation : x = 11 -y so it will be x = 11 - ...

OpenStudy (anonymous):

I get a fraction?

OpenStudy (anonymous):

14/5

OpenStudy (anonymous):

very good now plug it in

OpenStudy (anonymous):

Then i get x=41/5

OpenStudy (anonymous):

yes. so then you have x and y

OpenStudy (anonymous):

So whats y?

OpenStudy (anonymous):

we found it before...14/5